OpenStudy (ckenn):

A circle is centered at (−3, 2) and has a radius of 2. Which of the following is the equation for this circle?

OpenStudy (ckenn):

1.) (x + 3)2 + (y − 2)2 = 2 2.) (x − 3)2 + (y + 2)2 = 4 3.) (x + 3)2 + (y − 2)2 = 4 4.) (x − 3)2 + (y + 2)2 = 2

OpenStudy (ckenn):

@IrishBoy123 @Michele_Laino

OpenStudy (irishboy123):

\((x-a)^2 + (y-b)^2 = r^2\) is circle of radius r, centred on (a,b)

OpenStudy (ckenn):

so the y would turn positive because of the negative 3?

OpenStudy (irishboy123):

no, it,s (-3,2) the x co-ordinate of the centre is negative.....

OpenStudy (ckenn):

(x-(-3)^2 + (y-2)^2 = r^2

OpenStudy (ryandane):

So Irish you getting that or what ;)

OpenStudy (irishboy123):

@ckenn plug in for r too and i think you are in business :-))

OpenStudy (ryandane):

I see you hit her on the dm tell her come through ;) @IrishBoy123

OpenStudy (ckenn):

thank you (: @IrishBoy123 and ew @ryandane

OpenStudy (ryandane):

@ckenn um bored lol